OpenStudy (anonymous):

Which fractions are greater than -7/15 (the whole fraction is - not just the 7) ? There are 2 answers that are right

OpenStudy (anonymous):

\[-\frac{ 4 }{ 9 } \]

OpenStudy (anonymous):

\[-\frac{ 6 }{ 11}\]

OpenStudy (anonymous):

\[-\frac{ 5 }{ 11}\]

OpenStudy (anonymous):

\[-\frac{ 1 }{ 2}\]

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

- 1/2?

OpenStudy (skullpatrol):

why? :-)

OpenStudy (anonymous):

beacause 7 in a little less then 1/2 of 15

OpenStudy (skullpatrol):

Good try. With questions like these it is best to draw a number line and use the idea of common denominators

OpenStudy (anonymous):

okay

OpenStudy (skullpatrol):